OpenStudy (anonymous):

What the answer be? 1/3 or -2/3 for the value of x in: 15x-12y=13 30x+9y=4 Not sure, because both are possible answers. But only of the is correct.

OpenStudy (freckles):

so solve for x ... you tried to set the y's up for elmination that is...multiply the first equation by 9 and the second equation by 12 then add the equations together

OpenStudy (freckles):

so you have the following equation to solve: \[(9 \cdot 15+12 \cdot 30)x+(9 \cdot -12+12 \cdot 9)y=(9 \cdot 13+12 \cdot 4)\]

OpenStudy (anonymous):

Thanks. :)